Determine the energy of a swinging pendulum bob at its highest position, at the middle position, and at its lowest position which will be considered the reference level. The length of the string is 1 m and the height of its highest position is 1.5 m. The mass m of the bob is 0.20 kg.



1
Expert's answer
2021-11-21T17:24:29-0500

1. "E_1=mgh=0.2\cdot9.8\cdot1.5=2.94\ (J)"


2. "E_2=mgh_1+mv^2\/2=mgh_1+mg(h-h_1)=mgh="


"=0.2\cdot9.8\cdot1.5=2.94\ (J)"


3. "E_3=mv^2/2=mgh=0.2\cdot9.8\cdot1.5=2.94\ (J)"